A carom board ( 4 ft × 4 ft square ) has the queen at the centre. The queen, hit by the striker moves to the front edge, rebounds and goes in the hole behind the striking line. Find the magnitude of the displacement of the quen (a) from the centre to the front edge, (b) from the front edge to the hole and (c) from the centre to the hole.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. In ∆ABC , tan θ = x/2 and in ∆DCE, tanθ = (2-x) / 4 tan θ = (x/2) = (2 - x)/4 = 4x ⇒ 4 – 2x = 4x ⇒ 6 x = 4 ⇒ x = 2/3 ft (a) In ∆ABC, AC = √(AB^2+BC^2 ) = 2/3 √10 ft (b) In ∆CDE, DE = 1 – (2/3) = 4/3 ft CD = 4 ft. So, CE = √(CD^2+DE^2 ) = 4/3 √10 ft (c) Here the displacement vector r ⃗ = 7 i ̂ + 4 j ̂ + 3k ̂ (b) the components of the displacement vector are 7 ft, 4 ft, and 3 ft.